The greater of two numbers is 8 more than 5 times the smaller. Their sum is 68. Find the numbers

Answer by joecbaseball(37)

Let the bigger number be x, and let the smaller number be y.
Then x = 8 bigger than 5 times the smaller, so x = 5y + 8, and…
x + y = 68
Substitute the value of x (5y + 8) into the x value of the second equation, and you get…
(5y + 8) + y = 68, which gives you 6y + 8 = 68.
Subtract 8 from each side to get 6y = 60, and divide both sides by 6 to get y = 10.
Now substitute this value of y into either equation, and solve for x.
This will give you an x value of 58, and those are your answers;
x= 58 and y = 10.
Good luck